what percent of 25 is 16

Answer by stanbon(75887)

what percent of 25 is 16
------
x/100 = 16/25
----
x = 100(16/25)
---
x = 4*16
----
x = 64
-----
Ans: 64%
===================
Cheers,
Stan H.
=============